Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold to a Canadian importer for $904. Selected data for the company’s operations last year follow:

  Units in beginning inventory 0
  Units produced 23,000
  Units sold 20,000
  Units in ending inventory 3,000
  Variable costs per unit:
       Direct materials $ 160
       Direct labour $ 510
       Variable manufacturing overhead $ 60
       Variable selling and administrative $ 17
  Fixed costs:
       Fixed manufacturing overhead $ 790,000
       Fixed selling and administrative $ 860,000

Required:

1. Assume that the company uses absorption costing. Compute the unit product cost for one gamelan.

2. Assume that the company uses variable costing. Compute the unit product cost for one gamelan.
Answer #1

1. By Absorption Costing:-

Total per unit variable cost= 160+510+60+17=747

Total Per unit Fixed cost= (790000+860000/23000)=71.74

Total cost per unit =818.74

2. By Variable costing method:-

Total per unit variable cost= 160+510+60+17=747